ADTC meeting
Francois showed the follow-up of the YETS activities. Changes mentioned for the demi water cut (now 5-7th Feb) and the cryo plant maintenance (09 - 24th March). Crane refurbishment going on as planned, some limited cranage operations possible as of 12th Feb, under the responsibility of the contractor. To be limited to strict minimum as it could have an impact on the end of works date.
ALPHA barrack replacement: installation of the cables in the false floor is delayed till 12th Feb. No consequence.
BHN46 and QFN32 magnets are back in the ring and shielding is closed. 1 week ahead of time, but will not translate into early restart of AD machine.
C02 cavity power test delayed till week 7. No impact. Visits on ELENA viewing platform can resume.
ELENA ISTs complete. HW tests starting. Machine will restart with Hminus when demi water is back (possibly 8th Feb). Access will be needed 12-14Feb for LNE03 SEM installation.
Barabara reported for BASE activities during YETS. pbars were trapped on 27th october last year, and 21 pbars are still in there, without any loss since 15th December (showing vacuum is better than 10E-18 mbar !!!). Ultra precise measurements can only be done when the environment is quiet (Xmas break, WE, night). The ongoing measurements are expected to yield results with great improvement wrt earlier measurements. It is therefore of outmost importance that the trap can be powered and stay cold at all time (short possible power cuts announced in advance, lHe delivery continued at a pace of 1 dewar per week), and quiet times are preserved as often as possible, including during commissioning of the machines. The operation of GBAR with Hminus beyond the physics run may conflict with this requirement
Christian showed the status of BASE-STEP. The trap was removed from the zone after a principle transport test, to perform repair and improvement operations. The objective is to be ready for early March to be re-installed in the zone for cooldown. Consumption of lHe estimated to 30l/week. Then attempts will be done to improve the vacuum at the interface with ELENA transfer line. Protons transport will be tried, and when pbar trapping is demonstrated, transport attempts can be undertaken.
Laurette showed the machines restart planning. The AD target aims at DSO tests for 23rd Feb. First nominal PS beams on target expected for the 14th of March. HW commissioning between 4th and 14th of March, commissioning with beam as of 14th of March. Start of physics planned for 22nd April, maybe on 18th is all goes as planned... The end of physics run on 28th October. GBAR requested the possibility to run with ELENA Hminus till 20th december. Being evaluated.